OpenStudy (anonymous):

Does anyone get this? Change 4x - 3y = -15 to slope-intercept form and state the slope and y-intercept. You must show all work

OpenStudy (anonymous):

changing to slope intercept is achieved by isolating y in this case we subtract 4x from both sides and get -3y = -15 -4x then divide both sides by -3 .. what do you get ?

OpenStudy (anonymous):

-3 and -3 cancels -15/-3 = 5

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

y = (4/3)x + 5

OpenStudy (anonymous):

im wrong?

OpenStudy (anonymous):

this is what you should get after the division by -3

OpenStudy (anonymous):

oh, i made a few errors

OpenStudy (anonymous):

what's the outcome?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if you have y= mx + b m is the slope b is the y intercept so y = (4/3)x + 5 4/3 is the slope and 5 is the y intercept

OpenStudy (anonymous):

i solved it without seeing that!!!!! :)

OpenStudy (anonymous):

well im glad :)

OpenStudy (anonymous):

i have to go now .. see you later bye :)

OpenStudy (anonymous):

bye!!!!